OpenStudy (goformit100):

Ayush was born two years after his father's marriage. His mother is five years younger than his father but 20 years older than Ayush who is 10 years old. At what age did the father get married ? A. 23 years B. 25 years C. 33 years D. 35 years

OpenStudy (goformit100):

@Noemi95

OpenStudy (anonymous):

A. 23 years Ayush's present age= 10 years His mothers present age= (10+20)=30 years His fathers present age= (30+5)=35 years His father at the time Ayush was born= (35-10)= 25 years therefore, Ayush's father's age at the time of marriage= (25-2)= 23 years